3 4 x P(X=x) 0 .02 1 .08 2 .20 .40 .30 Given a student had unprotected sex more than 2 times, what is the probability that thplease explain how to do this
Answer #1

P(having unprotected more than 2 times)=P(X>2) =P(X=3)+P(X=4)=0.4+0.3 =0.7

P(number is fewer than 4 and more than 2) =P(X=3)=0.4

therefore P( fewer than 4 and more than 2 given more than 2 times)=P(X<4|X>2) =P(2<X<4)/P(X>2)=0.4/0.7=0.57

option C is correct